# Multi–charged geometries with cosmological constant

### v1

@@ 4 Charged branes with cosmological constant
In section 2 we constructed multi-charged black holes with nontrivial potentials for scalar fields, while in section 3 we constructed multi–charged black branes without such potential. This brings up a natural question: can one find brane counterparts of the solution from section 2? In this section we briefly review some previous work suggesting that existence of clean analytical expressions for such brane solutions is unlikely and discuss some interesting properties of the differential governing these brane configuration.
@@
To summarize, in this section we have analyzed solutions describing charged $p$–brane in the presence of the cosmological constant or, more generally, a nontrivial potential for scalar fields, such as ones discussed in section 2. Although unfortunately explicit analytical solutions seem to be beyond reach, we argued that the various expansions used in [18] to demonstrate existence and uniqueness of neutral geometries with regular horizons extend to the charged case as well. In the latter case, such solutions are fully specified by the mass and the set of charges. In the neutral case, there was also a unique regular solution that had the symmetries of the brane ansatz and AdSd asymptotics, and properties of this geometry were governed by a single first order ODE of the Abel type. In this section we found the extension of the Abel equation to the charged case, (4.9)–(4), whose solutions interpolate between the near horizon limit of the extremal branes (4.11) and the AdSd space at infinity. The boundary condition (4.12) uniquely specifies the solution.
@@ 5 Discussion
In this article we explored geometries produced by multi–charged black holes and black branes with and without the cosmological constant. Getting inspiration from known solutions of gauged supergravities, we began with an ansatz for black holes charged under vector and scalar fields, and we found a unique potential for the scalars that can lead to solutions of this type. Our potential generalizes the results from all known gauged supergravities to an arbitrary number of scalar fields in an arbitrary number of dimensions. We also constructed geometries sourced by branes with an arbitrary number of charges and analyzed their properties, including their relations with higher dimensional wormholes.
@@
In section 3 we extended the multi–charged solutions from black holes to branes and their intersections, and in the absence of the cosmological parameter, the final results are given by equations (3.1) and (3.3). These answers generalize all known brane intersections from string and M theory that follow the harmonic rules, but in contrast to those cases, where the relative co–dimensions of branes are always divisible by four, expressions (3.3) don’t have this restriction. As we showed in section 3.3, the restriction on co–dimension originates from very specific string theoretic couplings of the dilaton with fluxes, and once these couplings are relaxed, one can have more general intersections. Similarly, all string theoretic branes are fully specified by their charges under gauge fields, and scalars are fully determined by these parameters. In section 3.4 we extend this construction to uncorrelated gauge and scalar charges and construct the most general geometry (3)–(3) consistent with symmetries of the branes. As expected from the general analysis of [9, 1], additional uncorrelated charges lead to singular horizons, but in the case of four dimensional dilatonic black holes such singular solutions have found some interesting applications [52, 53], and we extended these most general local solutions to $p$ branes in arbitrary dimensions.
In the presence of the cosmological parameter, the analytical solutions for branes seems to be beyond reach, but in sections 3.2 and 4 we analyzed some properties of the relevant geometries, including their existence and uniqueness, using various expansions and arguments inspired by renormalization group flows developed in [18, 19]. In particular, in section 4 we reduced the problem of constructing the extremal solution to one first order Abel differential equation (4.9), and in section 3.2 we performed some expansions of brane solutions in the near horizon limit and closely related wormhole geometries.
@@ Appendix A Static black holes in gauged supergravities
In this appendix we briefly summarize some well–known solutions of gauged supergravities, which motivate our discussion of static black holes with a cosmological parameter presented in section 2. In every case we begin with the full action of gauged supergravity, and then discuss its truncation for known static configurations and quote the black holes solution. We also quote the scalar potentials which are reproduced by our general answer (2.25).
@@
We begin with the four–dimensional gauged supergravity introduced in [27]. The full action is given by [^16] In this appendix we denote the square root of the determinant of the metric by $e$ since this is common in the literature on gauged supergravity. Symbol $g$ is reserved for the cosmological parameter.
@@
$S=\int ed^{4}x\left[R-\frac{1}{2}(\partial{\vec{\phi}})^{2}+8{\bar{g}}^{2}\sum_{k=1}^{3}\cosh\phi_{k}-\frac{1}{4}\sum_{i1}^{4}e^{{\vec{a}}_{I}\cdot{\vec{\phi}}}(F_{(I)})^{2}\right]$ (A.2)
@@
The spherically-symmetric black hole solution for this action was found in [31], and it reads
@@
The action of the five–dimensional gauged supergravity is given by [28]
@@
This solution is reproduced as a special case of our general result (2.2), (2.25). Rotating extensions of the solution (A) were constructed in [34, 41].
@@
$S$ $=$ $\int\left[\star R-\frac{1}{2}\star d\phi\wedge d\phi-\frac{1}{X^{2}}\star F_{(2)}\wedge F_{(2)}+g^{2}\left(9X^{2}+\frac{12}{X^{2}}-\frac{1}{X^{6}}\right)\right].$ (A.12)
@@
The action of the seven–dimensional gauged supergravity is given by [30] [^17] We use notation of [42].
@@
Article [45] considered an interesting system that extends gauged supergravities to all dimensions while keeping only one gauge fields. The action was postulated to be
$L=R-\frac{(d-1)(d-2)}{4}(\partial\phi)^{2}-\frac{1}{4}e^{-(d-1)\phi}F^{2}+g^{2}(d-1)\left[(d-3)e^{\phi}+e^{-(d-3)\phi}\right]\,,$ (A.18)
and a charged version of the Kerr-AdS solution in arbitrary dimensions [10] was constructed. The static version of the solution is given by [45]:
@@ Appendix B Equations of motion for multi–charged black holes
$E_{\mu\nu}=R_{\mu\nu}-w_{1}\partial_{\mu}\psi\partial_{\nu}\psi-w_{2}(\sum\partial_{\mu}\psi_{I}\partial_{\nu}\psi_{I})+\sum_{I}e^{\gamma\psi_{I}}\left[F^{I}_{\mu\alpha}F_{\nu}^{I\alpha}-\frac{1}{2(d-2)}g_{\mu\nu}F^{I}_{\alpha\beta}F_{I}^{\alpha\beta}\right],$
@@
$E_{rr}$ $=$ ${\dot{A}}{\dot{B}}-{\dot{A}}^{2}-{\ddot{A}}-\frac{n{\ddot{C}}}{2C}+\frac{n{\dot{B}}{\dot{C}}}{2C}+\frac{n{\dot{C}}^{2}}{4C^{2}}-w_{3}{\dot{\psi}}^{2}-w_{4}\sum{\dot{\psi}}_{I}^{2}+\frac{2(n_{-})^{3}e^{2B}}{nC^{n}}\sum q_{I}^{2}e^{-\gamma\psi_{I}}\,,$
$E_{\theta\theta}$ $=$ $n_{-}+\frac{e^{-2B}}{2}\left[({\dot{B}}-{\dot{A}}){\dot{C}}-\frac{(n-2){\dot{C}}^{2}}{2C}-{\ddot{C}}\right]-\frac{2(n_{-})^{2}}{nC^{n-1}}\sum q_{I}^{2}e^{-\gamma\psi_{I}}\,,$ (B.2)
$E_{tt}$ $=$ $e^{2A-2B}\left[{\ddot{A}}+{\dot{A}}^{2}-{\dot{A}}{\dot{B}}+\frac{n{\dot{A}}{\dot{C}}}{2C}\right]-\frac{2(n_{-})^{3}e^{2A}}{nC^{n}}\sum q_{I}^{2}e^{-\gamma\psi_{I}}\,.$
@@
    Since $q_{I}$ should depend only on $s_{I}$, the expressions above must vanish. Setting $\nu_{3}=0$ is not a good option since it leads to $q_{1}=q_{2}=0$. Setting the square brackets to zero, we find [^19] For $k=2$, when both square brackets in equations (4) are the same, so vanishing of the relations (4) determines only $\nu_{4}$. In this case, the values of $\nu_{4}$ is determined by the requirement for $s_{2}$ to be absent in $q_{1}^{2}$ beyond the limit (4).
@@
    Expanding the equation $E_{rr}=0$ in powers of $M$, we find two constraints on parameters $(w_{3},w_{4})$, and once these constraints are imposed, all Einstein’s equations are satisfied. Collecting expressions for all relevant parameters, we find
@@
    $w_{3}=-\frac{nn_{-}(\nu_{3}\gamma)^{2}}{8[kn_{-}\nu_{3}-2]},\quad w_{4}=\frac{n\nu_{3}\gamma^{2}}{8}\,.$ (B.13)
@@
$\mu=-\frac{\gamma}{2},\quad\nu_{1}=-(n-1)\nu,\quad\nu_{2}=\nu_{3}=\nu,\quad\nu_{4}=-\frac{\nu_{1}}{2},\quad\nu_{5}=n\nu\,,$
$w_{1}=-\frac{\gamma^{2}n(n-1)\nu^{2}}{8[kn_{-}\nu-2]},\quad w_{2}=\frac{1}{8}n\gamma^{2}\nu,\quad q_{I}=2\sqrt{n\nu}Ms_{I}c_{I}\,.$ (B.17)
@@
The potential (B.16) was written in terms of harmonic functions $H_{I}$, and it terms of scalar fields if becomes
@@
We will now verify that once the relations (B) and (B.19) are imposed, equations for the vector and scalar fields are satisfied.
@@
$2\left[w_{2\gamma}\nabla^{2}\psi_{I}+w_{1\gamma}\nabla^{2}\psi\right]=\frac{2}{\gamma}\left[2w_{2\gamma}\nabla^{2}\ln H_{I}+w_{2\gamma}[-n_{-}\nu+(2-kn_{-}\nu)\frac{w_{1\gamma}}{w_{2\gamma}}]\nabla^{2}\ln H\right]$
@@
$\quad=\frac{n\gamma\nu}{2\sqrt{-g_{d}}}\left[-\frac{(2Ms_{I}c_{I}(d-3))^{2}}{H_{I}^{2}r^{d-2}}+g^{2}\partial_{r}\left[r^{d}H^{n\nu}\frac{\partial_{r}H_{I}}{H_{I}}\right]\right]\,.$ (B.23)
@@
The $g$–independent part of (B) matches that contribution from the two–form field strength to the right hand side of (B.22):
@@
$2\left[w_{2\gamma}\nabla^{2}\psi_{I}+w_{1\gamma}\nabla^{2}\psi\right]-\frac{\gamma}{2}e^{\gamma\psi_{I}}F_{I}^{2}=\frac{ng^{2}\gamma\nu}{2\sqrt{-g_{d}}}\partial_{r}\left[r^{d}H^{n\nu}\frac{\partial_{r}H_{I}}{H_{I}}\right]$
$\quad=\frac{n(n_{-}g)^{2}\gamma\nu r^{d-2}h_{I}H^{n\nu}}{2\sqrt{-g_{d}}}\left[n\nu\sum_{J}h_{J}+\frac{d-1}{n_{-}}-\frac{1}{H_{I}}\right]\,.$ (B.25)
@@
The scalar fields are constrained if $k\nu_{4}=1$, then (B) implies that
@@
Only $k-1$ scalars are independent. Variation of the action (2.20) with respect to these scalars gives a counterpart of equations (B)
@@ C.1 The neutral case
Now we recall that the ansatz preserves its form under reparameterizations of the radial direction, and this symmetry can be used to choose a convenient gauge. There is a generic situation, where $A-B$ is not constant and two special cases:
@@
    The first two two equations in (C) become equivalent, and they give
@@
    Them the last two equation in (C) decouple, and one arrives at the solution
@@
    We will not study this option further.
@@
Note that we have chosen convenient integration constants while going from $(A_{p},R_{p})$ to $(A,R)$: such constants can be adjusted by coordinate transformations, such as a rescaling of the time direction. The solution (C.1) has four parameters $(c_{0},c_{1},c_{2},a_{0})$ which are subject to the constraint (C.20). The parameter $c_{0}$ can be eliminated by shifting the scalar field.

### v2

@@ 4 Charged branes with cosmological constant
In section 2 we constructed multi-charged black holes with nontrivial potentials for scalar fields, while in section 3 we constructed multi–charged black branes without such potential. This brings up a natural question: can one find brane counterparts of the solution from section 2? In this section we briefly review some previous work suggesting that existence of clean analytical expressions for such brane solutions is unlikely and discuss some interesting properties of the differential equations governing these brane configuration.
@@
To summarize, in this section we have analyzed solutions describing charged $p$–branes in the presence of the cosmological constant or, more generally, nontrivial potentials for scalar fields, such as ones discussed in section 2. Although unfortunately explicit analytical solutions seem to be beyond reach, we argued that the various expansions used in [41] to demonstrate existence and uniqueness of neutral geometries with regular horizons extend to the charged case as well. In the latter case, such solutions are fully specified by the mass and the set of charges. In the neutral case, there was also a unique regular solution that had the symmetries of the brane ansatz and AdSd asymptotics, and properties of this geometry were governed by a single first order ODE of the Abel type. In this section we found the extension of the Abel equation to the charged case, (4.9)–(4), whose solutions interpolate between the near horizon limit of the extremal branes (4.11) and the AdSd space at infinity. The boundary condition (4.12) uniquely specifies the solution.
@@ 5 Discussion
In this article we explored geometries produced by multi–charged black holes and black branes with and without the cosmological constant. Getting inspiration from known solutions of gauged supergravities, we began with an ansatz for black holes charged under vector and scalar fields, and we found the unique potential for the scalars that can lead to solutions of this type. Our potential generalizes the results from all known gauged supergravities to an arbitrary number of scalar fields in an arbitrary number of dimensions. We also constructed geometries sourced by branes with an arbitrary number of charges and analyzed their properties, including their relations with higher dimensional wormholes.
@@
In section 3 we extended the multi–charged solutions from black holes to branes and their intersections, and in the absence of the cosmological parameter, the final results are given by equations (3.1) and (3.3). These answers generalize all known brane intersections from string and M theory that follow the harmonic rules, but in contrast to those cases, where the relative co–dimensions of branes are always divisible by four, expressions (3.3) do not have this restriction. As we showed in section 3.3, the restriction on co–dimension originates from very specific string theoretic couplings of the dilaton with fluxes, and once these couplings are relaxed, one can have more general intersections. Similarly, all string theoretic branes are fully specified by their charges under gauge fields, and scalars are fully determined by these parameters. In section 3.4 we extended this construction to uncorrelated gauge and scalar charges and constructed the most general geometry (3)–(3) consistent with symmetries of the branes. As expected from the general analysis of [24, 1], additional uncorrelated charges lead to singular horizons, but in the case of four dimensional dilatonic black holes such singular solutions have found some interesting applications [120, 121], and we extended these most general local solutions to $p$ branes in arbitrary dimensions.
In the presence of the cosmological parameter, the analytical solutions for branes seem to be beyond reach, but in sections 3.2 and 4 we analyzed some properties of the relevant geometries, including their existence and uniqueness, using various expansions and arguments inspired by renormalization group flows developed in [41, 42]. In particular, in section 4 we reduced the problem of constructing the extremal solution to one first order Abel differential equation (4.9), and in section 3.2 we performed some expansions of brane solutions in the near horizon limit and closely related wormhole geometries.
@@ Appendix A Static black holes in gauged supergravities
In this appendix we briefly summarize some well–known solutions of gauged supergravities, which motivate our discussion of static black holes with a cosmological parameter presented in section 2. In every case we begin with the full action of gauged supergravity, and then discuss its truncation for known static configurations and quote the black hole solution. We also quote the scalar potentials which are reproduced by our general answer (2.25).
@@
We begin with the four–dimensional gauged supergravity introduced in [78, 79]. The full action is given by [^16] In this appendix we denote the square root of the determinant of the metric by $e$ since this is common in the literature on gauged supergravity. Symbol $g$ is reserved for the cosmological parameter.
@@
$S=\int ed^{4}x\left[R-\frac{1}{2}(\partial{\vec{\phi}})^{2}+8{\bar{g}}^{2}\sum_{k=1}^{3}\cosh\phi_{k}-\frac{1}{4}\sum_{i=1}^{4}e^{{\vec{a}}_{I}\cdot{\vec{\phi}}}(F_{(I)})^{2}\right]$ (A.2)
@@
The spherically-symmetric black hole solution for this action was found in [85, 86], and it reads
@@
The action of the five–dimensional gauged supergravity is given by [80, 81]
@@
This solution is reproduced as a special case of our general result (2.2), (2.25). Rotating extensions of the solution (A) were constructed in [89, 90, 91, 92, 93, 101].
@@
$S$ $=$ $\int\left[\star R-\frac{1}{2}\star d\phi\wedge d\phi-\frac{1}{X^{2}}\star F_{(2)}\wedge F_{(2)}+g^{2}\star\left(9X^{2}+\frac{12}{X^{2}}-\frac{1}{X^{6}}\right)\right].$ (A.12)
@@
The action of the seven–dimensional gauged supergravity is given by [83, 84] [^17] We use notation of [102].
@@
Article [105] considered an interesting system that extends gauged supergravities to all dimensions while keeping only one gauge fields. The action was postulated to be [^18] To fit this expression in one line, we defined $\zeta=(d-1)(d-2)$.
$S=\int ed^{d}x\left[R-\frac{\zeta}{4}(\partial\phi)^{2}-\frac{1}{4}e^{-(d-1)\phi}F^{2}+g^{2}(d-1)\left[(d-3)e^{\phi}+e^{-(d-3)\phi}\right]\right],$ (A.18)
and a charged version of the Kerr-AdS solution in arbitrary dimensions [25, 26] was constructed. The static version of the solution is given by [105]:
@@ Appendix B Equations of motion for multi–charged black holes
$E_{\mu\nu}=R_{\mu\nu}-w_{1}\partial_{\mu}\psi\partial_{\nu}\psi-w_{2}(\sum\partial_{\mu}\psi_{I}\partial_{\nu}\psi_{I})+\sum_{I}\frac{e^{\gamma\psi_{I}}}{2}\left[F^{I}_{\mu\alpha}F_{\nu}^{I\alpha}-\frac{1}{2(d-2)}g_{\mu\nu}F^{I}_{\alpha\beta}F_{I}^{\alpha\beta}\right],$
@@
$E_{rr}$ $=$ ${\dot{A}}{\dot{B}}-{\dot{A}}^{2}-{\ddot{A}}-\frac{n{\ddot{C}}}{2C}+\frac{n{\dot{B}}{\dot{C}}}{2C}+\frac{n{\dot{C}}^{2}}{4C^{2}}-w_{1}{\dot{\psi}}^{2}-w_{2}\sum{\dot{\psi}}_{I}^{2}+\frac{(n_{-})^{3}e^{2B}}{2nC^{n}}\sum q_{I}^{2}e^{-\gamma\psi_{I}}\,,$
$E_{\theta\theta}$ $=$ $n_{-}+\frac{e^{-2B}}{2}\left[({\dot{B}}-{\dot{A}}){\dot{C}}-\frac{(n-2){\dot{C}}^{2}}{2C}-{\ddot{C}}\right]-\frac{(n_{-})^{2}}{2nC^{n-1}}\sum q_{I}^{2}e^{-\gamma\psi_{I}}\,,$ (B.2)
$E_{tt}$ $=$ $e^{2A-2B}\left[{\ddot{A}}+{\dot{A}}^{2}-{\dot{A}}{\dot{B}}+\frac{n{\dot{A}}{\dot{C}}}{2C}\right]-\frac{(n_{-})^{3}e^{2A}}{2nC^{n}}\sum q_{I}^{2}e^{-\gamma\psi_{I}}\,.$
@@
    Since $q_{I}$ should depend only on $s_{I}$, the expressions above must vanish. Setting $\nu_{3}=0$ is not a good option since it leads to $q_{1}=q_{2}=0$. Setting the square brackets to zero, we find [^20] For $k=2$, both square brackets in equations (4) are the same, so vanishing of the relations (4) determines only $\nu_{4}$. In this case, the value of $\mu$ is determined by the requirement for $s_{2}$ to be absent in $q_{1}^{2}$ beyond the limit (4).
@@
    Expanding the equation $E_{rr}=0$ in powers of $M$, we find two constraints on parameters $(w_{1},w_{2})$, and once these constraints are imposed, all Einstein’s equations are satisfied. Collecting expressions for all relevant parameters, we find
@@
    $w_{1}=-\frac{nn_{-}(\nu_{3}\gamma)^{2}}{8[kn_{-}\nu_{3}-2]},\quad w_{2}=\frac{n\nu_{3}\gamma^{2}}{8}\,.$ (B.13)
@@
$\mu=-\frac{\gamma}{2},\quad\nu_{1}=-(n-1)\nu,\quad\nu_{2}=\nu_{3}=\nu,\quad\nu_{4}=-\frac{\nu_{1}}{2},\quad\nu_{5}=n\nu\,,$ (B.17)
$w_{1}=-\frac{\gamma^{2}n(n-1)\nu^{2}}{8[kn_{-}\nu-2]},\quad w_{2}=\frac{1}{8}n\gamma^{2}\nu,\quad q_{I}=2\sqrt{n\nu}Ms_{I}c_{I}\,.$
@@
The potential (B.16) was written in terms of harmonic functions $H_{I}$, and it terms of scalar fields it becomes
@@
We will now verify that once the relations (B.17) and (B.19) are imposed, equations for the vector and scalar fields are satisfied.
@@
$2\left[w_{2\gamma}\nabla^{2}\psi_{I}+w_{1\gamma}\nabla^{2}\psi\right]=\frac{2}{\gamma}\left[2w_{2\gamma}\nabla^{2}\ln H_{I}+w_{2\gamma}[-n_{-}\nu+(2-kn_{-}\nu)\frac{w_{1\gamma}}{w_{2\gamma}}]\nabla^{2}\ln H\right]$ (B.23)
@@
$\quad=\frac{n\gamma\nu}{2\sqrt{-g_{d}}}\left[-\frac{(2Ms_{I}c_{I}(d-3))^{2}}{H_{I}^{2}r^{d-2}}+g^{2}\partial_{r}\left[r^{d}H^{n\nu}\frac{\partial_{r}H_{I}}{H_{I}}\right]\right]\,.$
@@
The $g$–independent part of (B.23) matches the contribution from the two–form field strength to the right hand side of (B.22):
@@
$2\left[w_{2\gamma}\nabla^{2}\psi_{I}+w_{1\gamma}\nabla^{2}\psi\right]-\frac{\gamma}{2}e^{\gamma\psi_{I}}F_{I}^{2}=\frac{ng^{2}\gamma\nu}{2\sqrt{-g_{d}}}\partial_{r}\left[r^{d}H^{n\nu}\frac{\partial_{r}H_{I}}{H_{I}}\right]$ (B.25)
$\quad=\frac{n(n_{-}g)^{2}\gamma\nu r^{d-2}h_{I}H^{n\nu}}{2\sqrt{-g_{d}}}\left[n\nu\sum_{J}h_{J}+\frac{d-1}{n_{-}}-\frac{1}{H_{I}}\right]\,.$
@@
The scalar fields are constrained if $k\nu_{4}=1$, then (B.17) implies that
@@
Only $k-1$ scalars are independent. Variation of the action (2.20) with respect to these scalars gives a counterpart of equations (B.25)
@@ C.1 The neutral case
Now we recall that the ansatz (C) preserves its form under reparameterizations of the radial direction, and this symmetry can be used to choose a convenient gauge. There is a generic situation, where $A-B$ is not constant and two special cases:
@@
    The first two equations in (C) become equivalent, and they give
@@
    Then the last two equations in (C) decouple, and one arrives at the solution
@@
    We will not discuss this degenerate option further.
@@
Note that we have chosen convenient integration constants while going from $(A_{p},R_{p})$ to $(A,R)$: such constants can be adjusted by coordinate transformations, such as rescaling of the time direction. The solution (C.1) has four parameters $(c_{0},c_{1},c_{2},a_{0})$ which are subject to the constraint (C.20). The parameter $c_{0}$ can be eliminated by shifting the scalar field.